# DocFinQA: A Long-Context Financial Reasoning Dataset

### 12 Jan 2024 v1

@@ Abstract
Research in quantitative reasoning within the financial domain indeed necessitates the use of realistic tasks and data, primarily because of the significant impact of decisions made in business and finance. Financial professionals often interact with documents hundreds of pages long, but most research datasets drastically reduce this context length. To address this, we introduce a long-document financial QA task. We augment 7,621 questions from the existing FinQA dataset with full-document context, extending the average context length for each question from under 700 words in FinQA to 123k words in DocFinQA. We conduct extensive experiments of retrieval-based QA pipelines and long-context language models on the augmented data. Our results show that DocFinQA provides challenges for even the strongest, state-of-the-art systems.
@@ 1 Introduction
Automated financial analysis is challenging because it involves reasoning over large volumes of textual and tabular data Azzi et al. (2019). Existing works on automating financial numerical reasoning focus on unrealistically specific document snippets Chen et al. (2021); Zhu et al. (2021). Datasets are often limited to narrowly focused document sections, failing to reflect the broader and more realistic scenarios faced by analysts Masson and Montariol (2020). Financial professionals typically sift through hundreds of pages per document, requiring deep understanding of both the content and structure to effectively navigate and extract pertinent information. Current long-document QA datasets such as Kočiský et al. (2018) do not test the quantitative reasoning skills needed in the financial domain.
In this work, we introduce DocFinQA, a new long-document financial question answering task. We extend the FinQA dataset of expert annotated questions and answers Chen et al. (2021) with full Securities and Exchange Commission (SEC) reports. This results in a significantly larger context in the DocFinQA dataset – by an average factor of 175 – than the context in the FinQA dataset. The resulting long-document QA task offers a more robust evaluation of a model’s reasoning capabilities over long contexts. In line with recent work on program synthesis for financial QA Koncel-Kedziorski et al. (2023), the questions in DocFinQA are annotated with Python programs to generate the answers, allowing for training and evaluating program synthesis models for use in realistic financial workflows.
Figure 1: DocFinQA extends FinQA to inputs with over 100K tokens. The resulting dataset is more challenging, causing GPT-4 to make nearly twice as many errors.
Using this new setup, we conduct detailed experiments of retrieval-based and long-context LLM systems. We study a typical retrieval pipeline consisting of chunking and encoding the document, searching for the best chunks given a question, and passing a pair of a question and top-$k$ chunks to a generative QA model Hsu et al. (2021). We also evaluate retrieval-free approaches using long-context LLMs Weston and Sukhbaatar (2023). Our results show that for LLMs to be effective in financial settings it is necessary to directly address the specific nuances of the financial domain, like large contexts; our dataset effectively tests these nuances.
@@ 2 Related Work
Prior studies in financial question answering focus on non-numerical reasoning Day and Lee (2016); Jørgensen et al. (2023); Maia et al. (2018). Short-context grounded numerical reasoning tasks were introduced with datasets such as FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021). Recently, understanding long documents has attracted more attention for tasks involving events Yang et al. (2018), table of contents Bentabet et al. (2020), and causal relations Mariko et al. (2022). However, to the best of our knowledge, ours is the first to address financial numerical QA by grounding it with extremely long documents – upwards of hundreds of pages of context for a given question.
Long document QA has been studied in NLP with the introduction of datasets such as SearchQA Dunn et al. (2017), NarrativeQA Kočiský et al. (2018), and QuALITY Pang et al. (2022). Due to the limited context size of LLMs, retrieval-based models are commonly used to filter irrelevant text Izacard et al. (2022); Lewis et al. (2020). Recently, advances in attention mechanisms Beltagy et al. (2020); Dao et al. (2022) and positional embeddings Press et al. (2021); Su et al. (2023) allow for conducting end-to-end grounded QA with context windows of more than 100k tokens. However, these methods suffer from loss of important context Zhang et al. (2023) and often fail to make full use of longer inputs Liu et al. (2023). Our work studies the intersection of numerical reasoning and long document processing, and our results show that there is still much room for improvement.
Dataset #Docs #QAs Multi-page Number Table #Words
@@
Table 1: Comparison of DocFinQA and existing datasets for Finance QA and Long Document QA.
@@ 3 DocFinQA Dataset
To assess the reasoning capabilities of LLMs for long-context financial question answering, we take the FinQA dataset as our starting point. FinQA consists of questions created by financial experts. These questions are grounded in text and tables that are extracted from the earnings reports of S&P 500 companies. Table 5 shows a QA data sample in FinQA.
The original FinQA dataset was constructed by financial professionals who knew the full document context for each question, but the released data only contains select content from a single page of the document. We extend this by providing the full SEC filing for each question, resulting in a more challenging task for modern LLMs. Additionally, the original task was to provide a numerical answer for each question. This provides no transparency for how the model arrived at its answer, which can limit user trust Ehsan et al. (2019); Lamm et al. (2020). We follow the setup of Koncel-Kedziorski et al. (2023) by requiring the model to generate executable Python code to produce an answer. This provides insight into how a model produced its answer – specifically, via illustrating what information from the context is used, along with what arithmetic operations are performed.
@@
Dataset Representation: Each question in FinQA is a triplet $(c^{golden},q,a)$ of a golden context $c^{golden}$, a question $q$, and an answer $a$ written in human language. We extend the dataset in two ways: (1) context $c^{golden}$ is extended to the full document context $D$; and (2) we added a Python program $p$ that produces the answer $a$. Each final sample in DocFinQA is a quartet $(D,q,p,a)$. An example of this can be seen in Appendix J.
Filings Collection: For each question of the FinQA dataset, we identify the corresponding SEC filing from which it was originally annotated. We retrieve the full filing in HTML/XML format from SEC EDGAR service [^1] https://www.sec.gov/edgar/. Figure 2 shows the distribution of document length in DocFinQA. Since each filing contains many tables, maintaining the structure and order during extraction is critical for the numerical reasoning task. We convert each HTML-formatted filing to PDF format and use Kensho Extract [^2] https://kensho.com/extract to parse the filing into text format with tables being represented in markdown format. This process ensures that: (i) our dataset is rooted in authentic and relevant financial documentation and (ii) all the tables in the filings are parsed with high precision into a consistent format without any HTML-tag noise. We discuss our parsing choices in more detail in Appendix D.
@@
Figure 2: Histogram of document length (#words) in DocFinQA dataset. Dash line shows the average length.
@@
Chunking and Alignment: To study retrieval-based QA systems, we split each document $D$ into a set of chunks $C=\{c_{1},\cdots,c_{n}\}$. Each chunk contains 2,750 characters (equivalent to approximately 509 tokens) with 20% overlapping to avoid missing important context at the edges. To compute the performance, from the chunk set $C$ associated with each document $D$, we identify the best context chunk, $c^{*}$ that includes the information to answer question $q$. Since FinQA already provides the $c^{golden}$, we compute pair-wise scores between the chunks in $C$ and the golden chunk $c^{golden}$. We found that four-gram based similarity score offers the sharpest matching signal among tri-gram, four-gram, and fuzzy matching. The chunk that yields the highest score is selected as the target context chunk for retrieval.
We verify that this process results in good $c^{*}$ chunks through manual inspection and by substituting $c^{*}$ for $c^{golden}$ in a few-shot QA evaluation. This evaluation using GPT-3.5 on the FinQA test set shows highly similar performance (67.5% vs 67.3% accuracy).
@@
Code Generation: The FinQA dataset provides solutions in a “program” syntax that, when executed, yields the golden result (e.g., subtract(34.8, 1.2), divide(#0, 34.8)). However, this derivation can be difficult to interpret and provides no meaningful context of what was being calculated. Koncel-Kedziorski et al. (2023) augments FinQA with readable Python code that can be executed to derive the answer, providing a layer of reliability and interpretability for the task at hand. Thus, we use the code-enhanced version of FinQA to construct DocFinQA (See Appendix F).
@@
Statistics: The resultant DocFinQA dataset comprises 5,798 training, 791 development, and 1,032 test samples, derived from a diverse set of 1,236 unique SEC filings. By encapsulating a broad temporal scope, our dataset not only addresses the immediate need for long-context financial data but also aligns with the evolving dynamics of the financial landscape. Table 1 shows the statistics and characteristics of DocFinQA in comparison with other finance numerical reasoning and long-document QA datasets.
@@
Due to the limited capability of the LLM we used to generate the code and the collection of the SEC filings, DocFinQA use a subset of FinQA questions. We analyze the distribution of the question types of FinQA and DocFinQA datasets (See Appendix A). We do not see any loss of any major question types in the dataset.
@@ 4 Retrieval-based QA Evaluation
Retrieval Task We test three encoding models for context retrieval: ColBERT (ColB.)Khattab and Zaharia (2020), Sentence-BERT (SentB.) Reimers and Gurevych (2019), and OpenAI’s Ada (Ada) Greene et al. (2022). Further, we finetune the ColBERT model (FT ColB.) on the training set of DocFinQA to measure performance using an in-domain model. We explored matching-based models (e.g., BM25 method Robertson et al. (1995)) but observe poor performance (See Appendix H).
@@
To retrieve context for a question $q$ over chunk set $C=\{c\}$, we encode both $q$ and $C$ with an encoding model. This results in a set of embeddings for the each chunk, $V(c_{i})$ and an embedding for the question, $V(q)$. We compute the cosine similarity between $V(c_{i})$ and $V(q)$ to retrieve the top-k most similar chunks, $\{r_{1},r_{2},..r_{k}\}$.

### 29 Feb 2024 v2

@@ Abstract
For large language models (LLMs) to be effective in the financial domain – where each decision can have a significant impact – it is necessary to investigate realistic tasks and data. Financial professionals often interact with documents that are hundreds of pages long, but most financial research datasets only deal with short excerpts from these documents. To address this, we introduce a long-document financial QA task. We augment 7,437 questions from the existing FinQA dataset with the full-document context, extending the average context length from under 700 words in FinQA to 123k words in DocFinQA. We conduct extensive experiments over retrieval-based QA pipelines and long-context language models. DocFinQA proves a significant challenge for even state-of-the-art systems. We also provide a case-study on the longest documents in DocFinQA and find that models particularly struggle on these documents. Addressing these challenges may have a wide reaching impact across applications where specificity and long-range contexts are critical, like gene sequences and legal document contract analysis. The data and code is publicly accessible at github.com/anonymous.
@@ 1 Introduction
The frequent need to reason over large volumes of textual and tabular data makes financial analysis particularly challenging for LLMs Azzi et al. (2019). Existing work on automating financial numerical reasoning focuses on unrealistically specific document snippets Chen et al. (2021); Zhu et al. (2021). Datasets are often limited to pre-selected document sections, failing to reflect the broader and more realistic scenarios faced by analysts Masson and Montariol (2020). Financial professionals often sift through hundreds of pages per document, requiring deep understanding of both the content and structure to effectively navigate and extract pertinent information. Current long-document QA datasets such as NarrativeQA Kočiský et al. (2018) do not test the quantitative reasoning skills needed in the financial domain.
In this work, we introduce DocFinQA, a long-document financial question answering task. We extend the FinQA dataset of expert annotated questions and answers Chen et al. (2021) with full Securities and Exchange Commission (SEC) reports. This results in a significantly longer context in the DocFinQA dataset – by a factor of 175 – than the FinQA dataset. The resulting long-document QA task offers a more realistic evaluation of a model’s reasoning capabilities over financial documents. In line with recent work on program synthesis for financial QA Koncel-Kedziorski et al. (2023), the questions in DocFinQA are annotated with Python programs to generate the answers, allowing for training and evaluating program synthesis models for use in realistic financial workflows.
Figure 1: DocFinQA extends FinQA to documents often over 150 pages long (100K+ tokens), so it is difficult to find the pertinent information. The question for the example above is: “For the quarter December 31, 2012 what was the percent of the total number of shares purchased in December?” The correct answer is 16.5%.
Using this setup, we evaluate retrieval-based and long-context LLM systems. We study a typical retrieval pipeline that chunks and encodes the document, searching for the best chunks given a question, and passing the question and top-$k$ chunks to a generative QA model Hsu et al. (2021).
We also evaluate retrieval-free approaches using long-context LLMs Weston and Sukhbaatar (2023). Our results show that the successful employment of LLMs in financial settings requires further study of the specific nuances of the financial domain, such as context disambiguation. Our dataset represents a step towards better capturing these nuances.
@@ 2 Related Work
Prior studies in financial question answering focus on non-numerical reasoning Day and Lee (2016); Jørgensen et al. (2023); Maia et al. (2018). Short-context grounded numerical reasoning tasks were introduced with datasets such as FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021). Recently, understanding long documents has attracted more attention for tasks involving events Yang et al. (2018), table of contents Bentabet et al. (2020), and causal relations Mariko et al. (2022). However, to the best of our knowledge, this is the first attempt to address financial numerical QA by grounded in long documents with upwards of hundreds of pages of context for each question.
Long-document QA has been studied in NLP with the introduction of datasets such as SearchQA Dunn et al. (2017), NarrativeQA Kočiský et al. (2018), QuALITY Pang et al. (2022), and PDFTriage Saad-Falcon et al. (2023). Due to the limited context size of LLMs, retrieval-based models are commonly used to filter irrelevant text Izacard et al. (2022); Lewis et al. (2020). Recently, advances in attention mechanisms Beltagy et al. (2020); Dao et al. (2022) and positional embeddings Press et al. (2021); Su et al. (2023) allow for end-to-end grounded QA with context windows of more than 100k tokens. However, these methods suffer from loss of important context Zhang et al. (2023) and often fail to make full use of longer inputs Liu et al. (2023). Our work studies the intersection of numerical reasoning and long-document processing, and our results demonstrate that there is still ample room for improvement in this domain.
Dataset #Docs #QAs #Words Pag Num Tab
@@
Table 1: Comparison of DocFinQA and existing Finance QA and Long Document QA dataset. DocFinQA includes multi-page documents rich with both numeric data and tables.
@@ 3 DocFinQA Dataset
Our dataset is an extension of the FinQA dataset. (See Table 5 for an example of FinQA). FinQA was created by finance experts from S&P 500 companies’ annual financial report (aka., 10-K). They used full documents when creating the questions (e.g., 100s of pages), but FinQA only includes the curated context necessary to produce correct answers (e.g., a paragraph and/or table of statistics).
To recreate the original problem, we retrieve the full SEC filing for each question. Following Koncel-Kedziorski et al. (2023), we require the model to generate Python code to produce an answer instead of directly generating the numeric answer. The generated Python code illustrates what information from the context is used, along with what arithmetic operations are performed.
@@
Dataset Representation: Each question in FinQA is a triplet $(c^{golden},q,a)$ composed of a golden context $c^{golden}$, a question $q$, and an answer $a$ written in human language. We extend the dataset in two ways: (1) context $c^{golden}$ is extended to the full document context $D$; and (2) we added a Python program $p$ that produces the answer $a$. Each final sample in DocFinQA is a quartet $(D,q,p,a)$. See Appendix M for an example.
Filings Collection: For each question of the FinQA dataset, we identify the corresponding SEC filing from which it was originally created. We retrieved the filing in HTML/XML format from SEC’s EDGAR service and parsed text and table into clean markdown format text Wang et al. (2023). The collection and parsing processes are presented in more detail in Appendix F and Appendix G, respectively. Figure 2 shows the distribution of document lengths in DocFinQA.
@@
Figure 2: Histogram of document length (#words) in DocFinQA dataset with dash line representing the average length of the documents. Purple line depicts the proportion of documents where the question context is within the current number of words.
@@
Chunking and Alignment: To study retrieval-based QA systems, we split each document $D$ into a set of chunks $C=\{c_{1},\cdots,c_{n}\}$. Each chunk contains 2,750 characters ($\sim\!509$ tokens) with a 20% overlap to avoid missing context at the edges. To compute the performance, we identify the best context chunk, $c^{*}$, from the chunk set $C$ associated with each document $D$ that includes the information to answer question $q$. Since FinQA already provides $c^{golden}$, we compute a pair-wise score $(c_{i},c^{golden})$, for all chunks, $c_{i}\in C$, including the golden chunk. We found that four-gram based similarity score offers the sharpest matching signal among tri-gram, four-gram, and fuzzy matching. The chunk with the highest score is selected as the target context chunk for retrieval. We verify that this process results in good $c^{*}$ chunks through manual inspection and by substituting $c^{*}$ for $c^{golden}$ in a few-shot QA evaluation with GPT-3.5. The model scores slightly higher when using $c^{*}$ rather than $c^{golden}$ (67.5% vs 67.3% accuracy respectively), suggesting that little context is lost.
@@
Code Generation: The FinQA dataset provides solutions in a “program” syntax that, when executed, yields the answer (e.g., in Figure 1 the solution is divide(102400, 619314). However, this derivation does not provide meaningful context of what is being calculated. In our running example, 102400 is not semantically grounded to the document. Koncel-Kedziorski et al. (2023) augments FinQA with readable Python code (including named variables like, dec_shares = 102_400) that can be executed to derive the answer, providing a layer of interpretability. Thus, we use the code-enhanced version of FinQA to construct DocFinQA (See Appendix I).
@@
Statistics: The resultant DocFinQA dataset comprises 5,735 training, 780 development, and 922 test samples, derived from 801 unique SEC filings. Table 1 shows the statistics and characteristics of DocFinQA in comparison with other finance numerical reasoning and long-document QA datasets.
@@
Due to the limited availability of complete SEC filings (refer Appendix F) and imperfections in the code generation process, DocFinQA encompasses a subset of FinQA questions. We analyze the distribution of the question types of FinQA and DocFinQA datasets (See Appendix C). DocFinQA remains a representation of all major question types from the original FinQA dataset.
@@ 4 Retrieval-based QA Evaluation
Retrieval Task We test three models for context retrieval: ColBERT (ColB)Khattab and Zaharia (2020), Sentence-BERT (SentB) Reimers and Gurevych (2019), and OpenAI’s Ada Greene et al. (2022). Further, we finetune the ColBERT model (FT ColB) on the training set of DocFinQA to evaluate an in-domain model. We test a matching-based model, BM25 Robertson et al. (1995), but observe poor performance (see Appendix K).
@@
To retrieve context for a question $q$ over chunk set $C$, we encode both $q$ and $C$ with the encoding models mentioned above. This results in an embedding for the question $v_{q}$ and chunk embeddings $V_{C}=\{v_{c_{i}}|c_{i}\in C\}$. We compute the cosine similarity between $v_{q}$ and each vector in $V_{C}$ to retrieve the top-k most similar chunks. We evaluate these models using hit rate (HR@k) on the test set of DocFinQA using the target $c^{*}$. Results are shown in Figure 3. The FT ColB yields the highest HR, followed by ColB. FT ColB yields an average improvement of 91% HR over SentB and obtains a 0.35 (HR@1) and 0.55 (HR@3).
Question Answering Task We formulate the QA task as a few-shot in-context learning task Brown et al. (2020). For each in-context example, we only provide the gold chunk and the correct answer. For the actual query we provide $k$ chunks. A sample prompt template is given in Figure 10.
We evaluate Falcon Penedo et al. (2023), MPT MosaicML (2023), LLaMa 2 and CodeLlaMa Touvron et al. (2023), Mistral Jiang et al. (2023), GPT3.5 Brown et al. (2020); OpenAI (2023) models. We weren’t able to evaluate proprietary models such as GPT3 Brown et al. (2020), BloombergGPT Wu et al. (2023) due to their inaccessibility. We skipped models that were not finetuned for code generation such as PIXIU Xie et al. (2023) and FinGPT Yang et al. (2023) due to their poor performance. We also skipped models trained for other languages such as BBT-Fin Lu et al. (2023) and XuanYuan 2.0 Zhang and Yang (2023).
Table 2 reports the performance of 11 state-of-the art models. Larger models outperform smaller models (e.g., MPT 30B vs MPT 7B). Models trained on code yield higher accuracy than non-code models (e.g., CodeLlama vs Llama). Models with additional supervised finetuning (e.g., LLama 2/7B+SFT) and instruction tuning (e.g., GPT-3.5) are among the best examined. Notably, Mistral 7B outperforms several larger models, although it lags behind Llama 2/7B+SFT and GPT-3.5.
The FT ColB model is the best retrieval model in all but one setting. It yields a marginal but consistent improvement over the ColB, and a large improvement over SentB and Ada.
Model/Size ColB SentB ADA FT ColB
Falcon/7B   2.3   0.3   1.2   1.8
MPT/7B   4.6   2.7   3.8   4.8
MPT/30B 17.3 11.1 12.1 18.1
Llama 2/7B 13.5 8.9 11.1 13.5
Llama 2/13B 18.7 12.7 14.9 19.1
CodeLlama/7B 15.6 12.2 15.2 16.8
CodeLlama/13B 19.1 13.8 18.8 21.0
Mistral/7B 23.2 14.9 21.5 25.0
Llama 2/7B+SFT 32.9 24.8 34.3 36.1
GPT-3.5/- 41.6 33.8 36.4 42.6
Table 2: Performance on DocFinQA test set. For each row, the best performance among all retrieval models is in bold. The fewshot setting is selected based on the best performance on the dev set (See Appendix L).
5 Case Study w/ 100K+ Token Documents
Recent LLMs can handle context lengths of 128K tokens, but more than 40% of the documents in DocFinQA remain unanswerable even at this content length (see Figure 2). Here, we evaluate performance on a test subsample of 200111The test subsample is limited to 200 documents due to the monetary and temporal costs of human evaluation and GPT4. long documents, each of which has 100K or more tokens.
We explore two retrieval-free options. System 2 Attention (S2A) extracts relevant information from each 100K-token chunk of a document before answering the question using the combined extracted information as context  (Weston and Sukhbaatar, 2023). The Iterative method produces the output program iteratively as the LLM processes each 100k section of the document. A temporary answer program (initially “None”) is input with each section to the LLM. We also report the performance of the best retrieval-based model (Retrieval) based on the experiment in Section 4. The human evaluation we conducted on these 200 questions highlights the challenging nature of this dataset. [^2] Experienced but non-expert human participants. See Appendix A for details.
Model/Size + Method w/ Retrieval Test Subsample
Human No 41.0
Mistral/7B + Iterative No 11.5
Mistral/7B + S2A No 15.5
Mistral/7B + Retrieval Yes 20.0
GPT-4 + Iterative No 20.0
GPT-4 + S2A No 23.0
GPT-4 + Retrieval Yes 47.5
Table 3: Retrieval-free performance on a case-study of $100K+$ token documents.
Non-expert human performance on DocFinQA is lower than FinQA (41% versus 50.7%). This can be attributed to the difficulty of also finding the golden page, compared to the golden page being given in FinQA. Notably, the expert performance reported in FinQA is 91.2%.
Still, non-expert human performance is double that of retrieval-free GPT-4 on these long documents, and roughly triple that of retrieval-free Mistral. Secondly, Iterative performed worse than S2A for both GPT-4 and Mistral with a reduced accuracy of 3% and 4%, respectively. Third, with retrieval, both Mistral and GPT-4 outperform their retrieval-free counterparts, with the assisted GPT-4 now on par with the human cohort. Together, these results highlight that DocFinQA is an effective and difficult test for long-document QA, and that there is still room for significant improvement in this domain. For instance, further exploration into methods that combine information across multiple calls to a document-processing LLM is warranted.
6 Conclusion
This paper introduces a realistic document-level question answering dataset over financial reports. Each question includes a full financial report (averaging 123K words), a far greater challenge than previous work that hones in on pre-specified content. Our findings reveal that this more realistic setting presents a significantly more difficult challenge, thereby opening new avenues for research in quantitative financial question answering.
7 Limitation
This work introduced an extension of the existing FinQA dataset. Due to limited human resources, we only validate the test set while the training and the development set were not fully validated. As a result, we can not make any claim of bias and question quality in the not-yet-validated data points offered in this paper. Additionally, as discussed in section 2, the code provided in this work was generated by WizardCoder LLMs. Our assumption is that the code is correct if it produces correct or approximately close to the golden answer. This method may generate both false positive codes (the code that generate correct answer with incorrect rationales) and false negative codes (the correct code that fail the approximation test).
8 Broader Impact and Ethical Considerations
We do not foresee any considerable risks associated with our work given that it is an extension of a publicly available documents and dataset. To uphold transparency, the paper provides detailed documentation of the dataset creation process, including the sources of data and annotation details. Our dataset serves as a resource to underscore the need for more long context oriented benchmarks both within and outside the financial domain and does not intend to criticize any one or more LLMs.
The annotation in this work is done automatically, so no crowd-source or contract annotators were hired throughout the process. The human evaluation in this study were done by full-time paid coworkers known to the authors.
Appendix A Human Evaluation Setting
We recruited three data professionals with 4-5 years of experience working with financial documents, including but not limited to 10-K filings, to estimate human evaluation. The professionals were provided with the entire document in PDF format, maintaining the SEC’s original format for ease of reading. They were allowed to use the keyword-search feature of PDF reader applications and a simple calculator for basic arithmetic operations required for this task. On average, the professionals spent 25 minutes per question.
Appendix B Retrieval Performance